# HoneyGPT: Breaking the Trilemma in Terminal Honeypots with Large Language Model

### v1

@@ 4.1.1. Concept of HoneyGPT
We define the interaction history as a sequence of N question-answer pairs:
(1) $\{(Q_{1},A_{1}),(Q_{2},A_{2}),\ldots,(Q_{N},A_{N})\}$
@@
To obtain the response $A_{i}$ from the $i$-th interaction, we constructed complex prompts. The operational logic of HoneyGPT during the $i$-th interaction is showcased in Figure 2. The $i$-th HoneyGPT interaction is formally defined by Equation  2.
(2) $(A_{i},C_{i},F_{i})=ChatGPT(P,S,H_{i},SR_{i},Q_{i})$
The components $P$, $S$, $H_{i}$, $SR_{i}$, and $Q_{i}$ together constitute the prompt for the $i$-th interaction. Here $P$ and $S$ represent the static configuration of the prompt, remaining unchanged throughout the interactions. $Q_{i}$ is the command sent by the attacker during the $i$-th interaction. $H_{i}$ and $SR_{i}$ are dynamic aspects of the prompt, updated after each interaction by the Prompt Manager M. The Prompt Manager synthesizes these variables into the new prompt for subsequent dialogue sessions, as defined by:
(3) $M(P,S,Q_{i},SR_{i},C_{i},F_{i},A_{i},w)=(P,S,H_{i+1},SR_{i+1},Q_{i+1})$
The following are definitions for each variable:
Honeypot Principle $P$ defines the strategies and guidelines used to create a honeypot.
Honeypot Setting $S$ refers to the initial configuration set within the prompt for the Honeypot.
History of Interaction $H_{i}$$\{(Q_{1},A_{1}),(Q_{2},A_{2}),\ldots,(Q_{i-1},A_{i-1})\}$ provides ChatGPT with the essential context required to support its capacity for extended interaction memory. To ensure that the length of $H_{i}$ remains within the context length limits of ChatGPT, the Prompt Manager $M$ implements a pruning algorithm to regulate the size of the historical context. This algorithm judiciously trims components of $H_{i}$ that have less significance, striking a balance between retaining pertinent context and maintaining the model’s operational efficiency within its architectural constraints.
System State Register $SR_{i}$ refers to a collection of system states communicated to ChatGPT at the $i$-th interaction: $\{(C_{1}),(C_{2}),\ldots,\\ (C_{i-1})\},$ where each state $C_{i}$ corresponds to the system’s changes after each prior interaction, paralleling the order of $H_{i}$. $SR_{i}$ is the manifestation of a ’chain-of-thought’ strategy, which HoneyGPT leverages to enhance its capacity for generating the subsequent response $A_{i}$.
Attacker Query $Q_{i}$ in the context of HoneyGPT, refers to the command issued by the attacker to the honeypot terminal system during the $i$-th interaction.
@@
Honeypot Answer $A_{i}$ in the context of HoneyGPT, denotes the command execution echo returned by the honeypot terminal to the attacker during the $i$-th interaction.
System State Evaluation $E_{i}$ is an assessment provided by ChatGPT regarding the system’s state post-interaction, consisting of two parts: Impact Factor $F_{i}$, which values the impact of $Q_{i}$ on the system, and $C_{i}$, which denotes the changes to the terminal system environment after $Q_{i}$ execution.
Weaken Factor $w$ is the constant used in the pruning algorithm, which accounts for the timing of interactions, preferentially trimming older and less relevant exchanges to maintain a recent, pertinent dialogue history.
4.1.2. Strategy
Figure 3. Comparison of the Effects of Integrating CoT
@@ 4.1.2. Strategy
The design of HoneyGPT incorporates a Chain of Thought (CoT) strategy to address the intrinsic limitations of ChatGPT in processing complex, extended dialogues and to circumvent the constraints of the prompt context length. The essence of CoT lies in compelling the LLM to decompose a multifaceted problem into a sequence of sub-problems, tackling them incrementally to enhance the LLM’s ability to comprehend and address intricate issues.
@@
Complex Extended Dialogue Tasks: In realistic adversarial scenarios, attackers often execute a series of malicious activities through multiple commands. As shown in Figure 3, attackers issue the ”write-elevate-execute” attack sequence. While native ChatGPT can effectively handle individual commands, it struggles with complex, interdependent, extended dialogue tasks that require memorization and coherence. By contrast, HoneyGPT can significantly increase the capability to handle such complex command sequences by incorporating the CoT strategy.
Within HoneyGPT, we simplify the complexity of extended dialogues into two steps: first, assessing the impact of previous commands on the operating system, and then determining the system’s response to new adversarial commands based on the evaluated impact. To ensure the continuity and automation of the attack interaction process, the task of assessing the impact of attack commands on the operating system has been delegated to ChatGPT. Thus, during each interaction, the prompt must provide ChatGPT with details of the changes that have occurred within the system up to that point( $SR_{i}$ ), as well as the present attack command( $Q_{i}$ ). ChatGPT is then tasked with analyzing the expected output from the terminal ( $A_{i}$ ) and determining the result of System State Evaluation ( $C_{i}$, $F_{i}$ ). Following the conclusion of an interaction, the Prompt Manager integrates $C_{i}$ into $SR_{i}$, respectively, setting the stage for the subsequent interaction.
Prompt Pruning Strategy: After each interaction, the Prompt Manager appends content to prompt, respectively, which results in an increasingly lengthy prompt for subsequent interactions. However, due to limitations imposed by computational resources, storage capacities, and model characteristics, LLMs have a restricted context length for prompts. Should the prompt’s token count exceed this threshold, ChatGPT will be unable to function. To address this issue, we introduce a strategy known as Prompt Pruning: if the prompt’s context length surpasses the limit, we eliminate the least impactful entries from the dynamic portion, $H_{i}$, with respect to the current operating system. Although $SR_{i}$ also grows dynamically, its length is significantly shorter than the context length limit, and therefore, it does not require pruning. Since $H_{i}$ represents the collection of historical honeypot interactions, it suffices to consider both the temporal sequence and the magnitude of impact to identify and remove the least pertinent interaction records for the current exchange. The impact magnitude is determined by ChatGPT during each interaction, for which we have provided a set of grading criteria that ChatGPT uses to generate $F_{i}$, as listed in Table 1.
4.2. Prompt Manager $M$
The operations of the Prompt Manager are categorized into three main processes: prompt construction, prompt pruning, and content updating.
Table 1. Rules of the Assignment of Values to $F_{i}$
@@ 0 Read file, display system information
1 file created, install tool
@@ 4 Impact services, delete files, password changed
4.2.1. Prompt Construction
The construction process takes place before each interaction session. When the attacker issues a command to HoneyGPT, the Prompt Manager formulates three critical questions for ChatGPT to answer, based on the Honeypot Principle, Honeypot Setting, Attacker Query, System State Register, and History of Interaction:
  1. (1)
    What is the terminal’s response ($A_{i}$) ?
  2. (2)
    How does the system state change ($C_{i}$) ?
  3. (3)
    Assign a numerical value to the command’s aggressiveness ($F_{i}$).
Question 2 underpins the CoT strategy, while Question 3 is intended for prompt length pruning and will be discussed later. Upon obtaining responses Ai, Ci, and Fi, HoneyGPT relays Ai back to the attacker, while Ci and Fi are utilized by the Prompt Manager to update the Hi and SRi in prompt for the subsequent interaction.
4.2.2. Prompt Pruning
Pruning is activated when the length of the prompt exceeds the operational limits of commercial LLMs. The primary objective is to trim interactions of less importance from the History of Interaction ($H_{i}$). Since the entries regarding system state changes are typically concise, $SR_{i}$ are exempt from this curtailment. The success of pruning depends on accurately singling out the least significant interactions from $H_{i}$. We have graded the impact of adversarial commands on the current system by using a scale from 0 to 4. The criteria for assigning values to the Impact Factor ($F_{i}$) are detailed in Table 1. LLMs will generate responses by referencing the table, ensuring that the pruning process is executed in a systematic and uniform manner.
In addition, the design contemplates the temporal dimension of command relevance, positing that earlier exchanges bear less pertinence to the present interaction. The Weaken Factor $w$ is introduced to quantify this temporal effect. With each interaction’s conclusion, the Impact Factor $F_{i}$ for each entry in $H_{i}$ is attenuated by multiplication with $w$, effectuating a time-based diminution of $F_{i}$. During the pruning process, entries with the lowest $F_{i}$ values are prioritized for removal to manage and preserve the prompt token limit efficiently.
Regarding the determination of the Weaken Factor $w$, both the length of the dialogue and the relevance of the content are considered. Observations from cybersecurity practices indicate that a sequence of attacks with strong correlation typically progresses through a ”write-elevate-execute” triad. Therefore, it is advisable to ensure that a newly executed command has a higher significance than a high-privilege command that occurred three steps prior, signified by $F_{i+3}\cdot w^{3}<F_{i}$. Additionally, the value of $w$ should not be too diminutive in order to adequately reflect the effect of time on the relevance of past commands. The process of temporal decay is implemented during Content Updating, where $F_{i}$ of each element is multiplied by the Weaken Factor ($w$).
4.2.3. Content Updating
The updating process is conducted at the end of each interaction session, aiming to renew the dynamic cache contents, specifically the System State Register $SR_{i}$ and the History of Interaction $H_{i}$. This does not include static elements such as constants S, P, and the adversary’s instructions $A_{i}$. Upon the completion of the $i$-th interaction, the Prompt Manager incorporates the query-response pair $(Q_{i},A_{i})$ into $H_{i}$, and the context $C_{i}$ into $SR_{i}$. The $SR_{i}$ supports the CoT strategy, while $H_{i}$ serves as contextual information for enriching ChatGPT’s understanding of the interaction history. Additionally, the updating includes the application of a temporal decay to the Impact Factor $F_{i}$ to support the pruning mechanism.
4.3. Configuration of HoneyGPT
Here we introduce the static components involved in configuring HoneyGPT, namely the System Principles ($P$) and Honeypot Settings ($S$), which remain unchanged once a honeypot is initialized. System Principles provide behavioral guidelines for simulating system activities, covering aspects such as the role of HoneyGPT, time sensitivity, input/output formats, and few-shot learning. These principles enable HoneyGPT to effectively attract and analyze potential attack behaviors by maintaining a credible system illusion. The Honeypot Settings involve specific configurations like hardware specifications and software environments, which are essential for simulating realistic targets that are attractive to attackers. These settings are designed to enhance the honeypot’s attractiveness and precision in capturing malicious exploitation such as cryptocurrency mining. Once set, both components provide a stable platform for HoneyGPT to operate effectively within its designed parameters. A detailed description of these two sets is given in Appendix A.
5. Evaluation
@@
Our evaluation of HoneyGPT includes two parts: the baseline evaluation and the field evaluation. The baseline evaluation compares the capabilities in deception, interaction level, and flexibility between HoneyGPT and traditional honeypots. Specifically, to evaluate the honeypots’ capability in deception and interaction level under consistent attack scenarios, the baseline evaluation utilizes the attack dataset collected by Cowrie. The field evaluation involves deploying HoneyGPT and Cowrie simultaneously on the Internet for a period of four weeks to assess whether HoneyGPT can influence attackers’ attack strategies and capture a wider range of attack types and interaction lengths.
5.1. Baseline Comparison Set
@@
Table 2 shows the experimental setup for the baseline comparison. The comparison of deception is conducted among HoneyGPT based on GPT-3.5 turbo interface, HoneyGPT based on GPT-4 turbo interface, Cowrie, and a real system. The interaction level comparison is limited to the comparisons among HoneyGPT based on GPT-3.5 turbo interface, HoneyGPT based on GPT-4 turbo interface, and Cowrie, because the real system exhibits the best interaction level and there is no need to compare. The flexibility comparison is conducted among Honeyd, a low-interaction honeypot, along with HoneyGPT, Cowrie, and the real system.
To assess the differences in deception and interaction level between HoneyGPT and traditional honeypots under the same attack scenarios, we use the same attack dataset to replay traffic, simulating identical attack scenarios. This dataset compiles attack data previously captured by Cowrie, a mainstream, code-based, medium-interaction open-source honeypot system  (Priya and Chakkaravarthy, 2023; Başer et al., 2021). As shown in Table 3, these datasets include a substantial volume of attack commands and valid attack sessions. A sequence of actions performed by an attacker post-successful login is considered as a valid attack session. We perform data cleaning and aggregation on these datasets, resulting in 160 attack sessions and 1,489 attack commands, which are used for baseline comparison.
Table 2. Baseline Comparison Set
Capability Type Subjects Method Research Type
Deception HoneyGPT (GPT-3.5 turbo), HoneyGPT (GPT-4), Cowrie, Real System Replay attack dataset Quantitative Research
Flexibility HoneyGPT, Cowrie, Honeyd, Real System Verification Individually Qualitative Research
Table 3. Attack Data Source
@@
 (Başer et al., 2021) Cowrie captured 2021/04/06-2021/06/11 5099153 2479
 (Priya and Chakkaravarthy, 2023) Cowrie captured 2022/06/09-2022/10/29 1261689 7374
5.2. Baseline Comparison of Deception

### v2

@@ IV-A The Constitution of Prompt
We first define ChatGPT’s interaction history as a sequence of N question-answer pairs:
$\{(Q_{1},A_{1}),(Q_{2},A_{2}),\ldots,(Q_{N},A_{N})\}$ (1)
@@
To obtain the response $A_{i}$ from the $i$-th interaction, we construct complex prompts. The $i$-th HoneyGPT interaction is formally defined by Equation  2.
$(A_{i},C_{i},F_{i})=ChatGPT(P,S,Q_{i},SR_{i},H_{i})$ (2)
As shown in Figure 1, the constitution of the prompt for the $i$-th interaction encompasses six components: Command $Q_{i}$ , Question Enhancement, Honeypot Principle $P$, Honeypot Setting $S$, System State Register $SR_{i}$ , and History of Interaction $H_{i}$. Here $P$ and $S$ represent the static configuration of the prompt, remaining unchanged throughout the interactions. $SR_{i}$ and $H_{i}$ are dynamic memory aspects of the prompt, which are updated after each interaction by the Prompt Manager. The definitions of variables used in HoneyGPT are given below.
  • Honeypot Principle $P$ defines the Principle that the LLMs should follow when performing honeypot tasks
  • Honeypot Setting $S$ refers to the Honeypot initial configuration set.
  • Attacker Query $Q_{i}$ denotes the command issued by the attacker during the $i$-th interaction.
@@
  • Honeypot Answer $A_{i}$ denotes the terminal output returned to the attacker during the $i$-th interaction.
  • History of Interaction $H_{i}$, represents the record of interactions between attackers and the honeypot prior to the i-th interaction, is defined as follows: $\{(Q_{1},A_{1}),(Q_{2},A_{2}),\ldots,(Q_{i-1},A_{i-1})\}$.
  • System State Register $SR_{i}$ refers to a collection of system states communicated to LLMs at the $i$-th interaction: $\{(C_{1}),(C_{2}),\ldots,(C_{i-1})\},$ where each state $C_{i}$ corresponds to the system’s changes after each prior interaction, paralleling the order of $H_{i}$.
  • New Change $C_{i}$ denotes the changes to the system environment after $i$-th interaction.
  • Impact Factor $F_{i}$ quantifies the impact of $Q_{i}$ on the system.
  • Weaken Factor $w$, utilized in the pruning algorithm, systematically diminishes the efficacy of $Q_{i}$ by scaling it with $w$ after each interaction.
IV-B The Strategy of Prompt Manager
@@
The design of HoneyGPT introduces strategies based on the Chain of Thought (CoT) for question enhancement and memory pruning. These strategies are developed to address the intrinsic limitations of ChatGPT in processing complex, extended dialogues and to circumvent the constraints of the prompt context length. The essence of CoT lies in compelling the LLM to decompose a multifaceted problem into a sequence of sub-problems, tackling them incrementally to enhance the LLM’s ability to comprehend and address intricate issues.
IV-B1 Question Enhancement for Complex Extended Dialogue Tasks
@@
Adversarial scenarios often involve attackers executing a series of malicious commands, such as the ”write-elevate-execute” sequence depicted in Figure 3. Native ChatGPT, while adept with individual commands, struggles with complex, interdependent dialogues requiring coherence and memory. In contrast, HoneyGPT’s integration of a Question Enhancement strategy markedly improves its capacity to process these complex sequences. Within HoneyGPT, we enhance the rudimentary long-dialogue response tasks of the honeypot into three specific sub-tasks:
  • Given the system state, what is the terminal’s output ($A_{i}$) ?
  • How does the system state change ($C_{i}$) ?
  • Assign a numerical value to the command’s effect ($F_{i}$).
To ensure the continuity and automation of the attack interaction process, the tasks of evaluating the impact of attack commands on the operating system, namely questions 2 and 3, have been delegated to ChatGPT. The analysis results from these tasks provide support for question 1 by supplying system state information in subsequent sessions. The question enhancement mechanism refines the questions and leverages ChatGPT to analyze system changes, thereby enhancing its consistency and effectiveness in handling complex, long-dialogue issues.
Figure 3: Comparison of the Effects of Question Enhancement
IV-B2 Memory Pruning for Prompt Length Limitation
As interactions accumulate, the dynamic memory ($H_{i}$) increases, leading to a longer prompt for subsequent interactions. However, due to limitations imposed by computational resources, storage capacities, and the inherent characteristics of the self-attention mechanism, LLMs have a restricted context length for prompts. Exceeding this limit renders ChatGPT non-functional. To address this issue, we introduce Memory Pruning algorithm: when the context length of a prompt surpasses the predefined threshold, we prune the least impactful entries from $H_{i}$ to balance the consistency and usability of HoneyGPT. While the $SR_{i}$ also increases dynamically, its length remains well below the context length limit, thus obviating the need for its pruning. The impact of $H_{i}$’s each entry is assessed by ChatGPT using the quantitative criteria specified in Table I. Furthermore, to account for the diminishing influence of earlier commands on the Impact Factor ($F_{i}$) over time, we introduce a Weaken Factor ($w$) that is applied to each interaction’s $F_{i}$. When selecting the value of $w$, it is crucial to ensure that a newly executed command carries more weight than a high-privilege command that occurred three steps prior. This principle can be represented mathematically as $F_{i+3}\cdot w^{3}>F_{i}$. Moreover, the value of $w$ should not be excessively small, as it must adequately capture the effect of time on the relevance of past commands.
TABLE I: Rules of the Assignment of Values to $F_{i}$
@@ 0 Read file, display system information
1 Create file, install tool
@@ 4 Impact services, delete files, password changed
IV-C The Workflow of Prompt Manager
As shown in Figure 2, The Prompt Manager orchestrates the interaction cycle in HoneyGPT, systematically handling attack commands and constructing precise prompts. This section delineates the workflow, emphasizing the integration of dynamic memory management and question enhancement strategies. Given that the Terminal Protocol Proxy solely handles protocol parsing and encapsulation, its discussion is omitted here. The workflow of the Prompt Manager is as follows: (1) Receive Attack Command: The attacker sends an attack command, denoted as $Q_{i}$. (2) Construct Prompt: The Prompt Manager constructs a comprehensive prompt, integrating the Question Enhancement mechanism IV-B1, which decomposes the interaction task into three sub-tasks.Additionally, the prompt includes the attacker’s command $Q_{i}$, and combines static configuration elements such as $S$ and $P$ with dynamic memory components $SR_{i}$ and $H_{i}$. (3) Interaction with ChatGPT: The constructed prompt is dispatched to ChatGPT for processing. (4) Receive Response from ChatGPT: ChatGPT returns a response, which includes the terminal output $A_{i}$, system changes $C_{i}$, and impact factor $F_{i}$. (5) Memory Maintenance: is divided into Memory Updating and Memory Pruning. Memory Updating integrates the current session’s memory into the prompt’s dynamic memory. Specifically, it appends $(Q_{i},A_{i})$ to $H_{i}$ and adds $C_{i}$ to $SR_{i}$. Additionally, $F_{i}$ is stored in the list $FL$ to inform Memory Pruning. Memory Pruning manages the prompt’s length. It first applies time decay by multiplying each session record’s impact factor in $FL$ by the Weaken Factor $w$. Next, it checks if the prompt length is approaching its limit. If it is, it identifies the index $j$ of the smallest $F_{j}$ in $FL$ and removes $(Q_{j},A_{j})$ from $H_{i}$ and $F_{j}$ from $FL$ to maintain efficiency. (6) Response attacker: The refined response $A_{i}$ is delivered back to the attacker, concluding the interaction sequence with a well-informed, precise output.
IV-D Static Configuration of HoneyGPT
Here we introduce the static components involved in configuring HoneyGPT, namely the System Principles ($P$) and Honeypot Settings ($S$), which remain unchanged once a honeypot is initialized.
IV-D1 System Principles $P$
System Principles provide behavioral guidelines for simulating system activities, covering aspects such as the role of HoneyGPT, time sensitivity, input/output formats, and few-shot learning. These principles enable HoneyGPT to effectively attract and analyze potential attack behaviors by maintaining a credible system illusion. HoneyGPT, serving as a terminal honeypot system, is specifically designed to attract and monitor attack activities. To achieve this objective, we have implemented the system’s principles within the prompts, providing operational guidelines for ChatGPT. These prompts play different crucial roles in the functionality of the system, including:
Role of HoneyGPT HoneyGPT strategically simulates an authentic terminal environment, tailored to align closely with attackers’ intents, thereby encouraging deeper and more sustained engagement. The system expertly balances authenticity with the objective of enticing attackers into prolonged interactions, thereby enhancing the honeypot’s effectiveness.
Time Sensitivity In scenarios where attacks are time-sensitive, attackers may issue commands related to current time inquiries, such as uptime and top. Since ChatGPT inherently lacks the capability to query real-time networked time data, it is necessary for us to provide time-related information such as the current time and boot time in the prompts to ChatGPT. This approach ensures that the honeypot’s responses are temporally accurate, thus more convincingly mimicking a real system environment.
Input/Output Format The structure of both input and output prompts critically defines each interaction with the LLM in HoneyGPT. The input format is crucial for the LLM’s accurate understanding of the honeypot system’s state and the commands issued by attackers. Conversely, the output format affects HoneyGPT’s operational effectiveness. To ensure HoneyGPT’s robustness, we instruct the LLM to use the JSON format for outputs. This standardization ensures consistent and correct parsing of responses during each interaction with the model.
Few-Shot Learning HoneyGPT utilizes a few-shot learning strategy to enhance ChatGPT’s comprehension of task requirements and improve the quality of its outputs, compensating for gaps in several command knowledge. Our experience indicates that by integrating only 4-5 representative examples, HoneyGPT’s generative capabilities can be significantly enhanced.
IV-D2 Honeypot Setting $S$
The Honeypot Settings describe the terminal system information simulated by HoneyGPT, which is crucial for attracting and capturing potential attackers, involving specific configurations like hardware specifications and software environments. These settings are designed to enhance the honeypot’s attractiveness and precision in capturing malicious exploitation such as cryptocurrency mining. Our experience indicates that honeypots with high-end configurations are more attractive to malicious entities, such as cryptocurrency-mining malware. Furthermore, the more details of honeypot configurations, the higher precision of the responses. The ability to customize honeypot settings via natural language significantly enhances flexibility and reduces the complexity of deployment. Honeypot settings are divided into two main categories: hardware and software.
Hardware It is essential to define the hardware specifications that will be emulated by HoneyGPT. It includes defining CPU types and counts, GPU presence, and storage capacities, those attributes that attackers frequently scrutinize. The specifications selected must mirror those of the systems intended to be simulated by HoneyGPT to maintain the authenticity of the honeypot environment. Additionally, configuring systems with high-end GPUs and CPUs can attract specific types of threat actors, such as crypto-mining malware, prompting deeper interactions.
Software The software environment of a honeypot, including the operating system, open ports, user configurations, and application services, must be strategically crafted to reflect the behaviors of potential targets. For instance, if the objective is to mimic a web server, HoneyGPT should be configured with the appropriate web server software, along with associated services and open ports. Additionally, attackers are often drawn to details such as system processes, resource utilization, scheduled tasks, user configurations, and filesystem information. These elements can be specifically tailored within HoneyGPT to create customized responses for each engagement scenario.
V Evaluation
@@
Our evaluation of HoneyGPT includes two parts: the baseline evaluation and the field evaluation. The baseline evaluation compares the capabilities in deception, interaction level, and flexibility between HoneyGPT and traditional honeypots. Specifically, to evaluate the honeypots’ capability in deception and interaction level under consistent attack scenarios, the baseline evaluation utilizes the attack dataset collected by Cowrie. The field evaluation involves deploying HoneyGPT and Cowrie simultaneously on the Internet for a period of three months to assess whether HoneyGPT can influence attackers’ attack strategies and capture a wider range of attack types and interaction lengths.
V-A Baseline Comparison Set
TABLE II: Baseline Comparison Set
Capability Type Subjects Method Research Type
Deception HoneyGPT (GPT-3.5 turbo), HoneyGPT (GPT-4), Cowrie, real-system Replay attack dataset Quantitative Research
Flexibility HoneyGPT, Cowrie, Honeyd, real-system Verification Individually Qualitative Research
@@ V-A Baseline Comparison Set
Table II shows the experimental setup for the baseline comparison. The comparison of deception is conducted among HoneyGPT based on GPT-3.5 turbo interface, HoneyGPT based on GPT-4 turbo interface, Cowrie, and a real system. The interaction level comparison is limited to the comparisons among HoneyGPT based on GPT-3.5 turbo interface, HoneyGPT based on GPT-4 turbo interface, and Cowrie, because the real system exhibits the best interaction level and there is no need to compare. The flexibility comparison is conducted among Honeyd, a low-interaction honeypot, along with HoneyGPT, Cowrie, and the real system.
(a) SALC,SALNLC,FALC,FALNLC
(b) Accuracy,Temptation,Attack Success Rate,OS Logic Compliance
Figure 4: Deception assessment results of HoneyGPT
To assess the differences in deception and interaction level between HoneyGPT and traditional honeypots under the same attack scenarios, we use the same attack dataset to replay traffic, simulating identical attack scenarios. This dataset compiles attack data previously captured by Cowrie, a mainstream, code-based, medium-interaction open-source honeypot system  [23, 3].
TABLE III: Attack Data Source
@@
 [3] Cowrie captured 2021/04/06-2021/06/11 5099153 2479
 [23] Cowrie captured 2022/06/09-2022/10/29 1261689 7374
As shown in Table III, these datasets include a substantial volume of attack commands and valid attack sessions. A sequence of actions performed by an attacker post-successful login is considered as a valid attack session. We perform data cleaning and aggregation on these datasets, resulting in 160 attack sessions and 1,489 attack commands, which are used for baseline comparison.
V-B Baseline Comparison of Deception